OpenStudy (anonymous):

what is 9 4/10 as a decimal

jimthompson5910 (jim_thompson5910):

0.70 = 70/100 reduce this as much as possible to get your final answer

OpenStudy (anonymous):

thank u.

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

also 9 4/10 = 9+4/10 = ???